Status Report

NASA Space Station On-Orbit Status 24 August 2004

By SpaceRef Editor
August 24, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 24 August 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

After wakeup at their regular 2:00am EDT (6:00am GMT), followed by morning inspection and breakfast, the crew immediately started out on a very busy day of systems maintenance, for CDR Padalka taking on the KOB thermal control system, for FE/SO Fincke tackling EMU restoration.

Mike continued the repair of EMU/spacesuit #3013 by running tests on its cooling function after yesterday’s work, during which he cleaned the cup that houses the pump rotor and installed a new impeller/rotor delivered on 15P.  Today’s cooling system tests were successful, and it appears that a big step was done toward regaining a fully functional EMU.  Tomorrow’s task will be to replace the water pump filter, close up the backpack and clean up.   [After activation of the EMU in the U.S. Airlock, which required temporary inhibition of the 400MHz transmitter of the Russian/European GTS (Global Timing System), the first part of the testing was to measure the current (amperage) used with the pump running dry, i.e., without fluid flow.  Then the test was repeated but this time with fluid flowing through the pump.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Gennady Padalka meanwhile worked on the Service Module (SM)’s no. 1 & 2 internal cooling loops, deinstalling a failed SPN pump panel and replaced it with a spare of new design on TsUP/Moscow instructions.  In order for the ground to monitor the work, it required Gennady to frequently connect/disconnect cabling between the BSR Regul interface unit (part of the Russian radio control and communications system) and the BITS2-12 onboard data/telemetry system.  The IFM also required temporary deactivation of Elektron, SKV-1, BMP and Vozdukh (if in Auto mode).   [Each of the two thermal loops (KOB-1 & KOB-2) has two pump panels, each of which has two pumps hard welded to it.  The automatic command algorithm does not allow a panel to run on only one of them, thus failure of one pump effectively fails the entire panel.  The replacement today used a new design with individually replaceable pumps, so that a pump failure will not require swap-out of the entire SPN.  The other SPNs were replaced on 7/16.]

Mike Fincke had a total of four hours blocked out for upgrading SSC (station support computer) laptops with new post-15P software delivered on CD-ROM (DVD)s.   [The reloading involved the SSC File Server (FS) laptop and six SSC “Clients”, i.e., NGSD (new generation support disk) IBM A31p ThinkPad laptops.  The reloads installed SSC FS version 4.00 on the FS and SSC Client A31p version 8.00 on the NGSDs.  The job also included post-reload cleanup and administrative work such as re-establishing the “partnership” between the SSCs and the crew’s Hewlett-Packard PDA handheld for file transfers & e-mail synchronization, and checking out the new Hazmat (hazardous materials) database application v.5.0 for the A31p’s.]

The CDR terminated the latest run of the Molniya-SM/LSO (GFI-10) geophysical experiment, started on 8/21, by deactivating the EGE 1 computer, the LSO 06 instrument and the Russian RBS power outlet.  On 8/29, he will copy the accumulated data to HDD (hard disk drive) on 8/29.   [Objective of the often-repeated geophysical Molniya-SM, similar to the French LSO experiment, is to record storm phenomena and other related events in the Earth’s equatorial regions.  The payload uses the French-provided EGE-1 laptop running the latest NORAD orbital parameters (TLEs, two-line elements).]

Padalka serviced the Matryoshka radiation payload by clearing its folders in the Matryoshka server (BSPN) from the Russian “Wiener” laptop (using a program called “ShellForKE”) for new measurements from the radiation monitoring equipment.  The procedure had to be timed with Matryoshka deactivation between comm passes.

Gennady dismantled the V1/V2 fan assembly, which Alex Kaleri had removed from the Descent Module (SA) of Soyuz-213/7S on 4/13.  The functioning V1 fan was prepared for return to the ground, wrapped shockproof in “vogatur” fabric, the failed V2 fan was disposed of.   [The assembly had been replaced in the DM’s cooler/dryer unit (KhSA) in a major IFM after one of the two fans failed on 10/18/03 during 7S free flight, resulting in higher-than-expected humidity (18 mmHg) in the Soyuz cabin for Expedition 8 on their flight to the ISS (10/18-10/20/03).]

The CDR also performed the daily routine inspection of the SM’s SOZh life support system (including replacement of ASU toilet facility inserts) and then prepared the regular IMS (inventory management system) “delta” file for export/import to the IMS databases.

At specific times during the day, as time permitted, Padalka worked a new task-listed run of the Uragan (“hurricane”) earth-imaging program, using the Kodak 760 DSC (digital still camera) with 800mm-lens from SM windows.  [Today’s targets for his photo imagery were the Kerch Strait dam, logging sites on mountain slopes S of Krasnodar, glaciers on the N slope of the Caucasus Ridge, the Kolka glacier on Mt. Kazbek, the city of Tbilisi, Caspian Sea shores, Mt. Elbrus, coastlines of Southern Europe, Cyprus, Israel, panoramic views of the Pyrenees, Palma de Mallorca, and volcanoes in South America.]

Mike Fincke was provided with a new list of options for his next “Saturday Science” program for his choice.   [Suggested by POIC (Payload Operations & Integration Center) were CFE (Capillary Flow Experiment) tests, an ISSI (In-Space Soldering Investigation) test (#3 of 5), an MFMG (Miscible Fluids in Microgravity) test, an EPO demo of the Blue Harp music instrument and/or the Puzzles toy, AVI movie file management & downlinking for ISSI, and operations with the HEAT experiment (investigating heat transfer performance of heat pipes).]

TsUP/Moscow continued the current extensive data collection and testing of the Russian ASN (satellite navigation system).   [Today, the U.S. handed over ISS attitude control to the Russian segment (RS) at 1:30pm EDT, for maneuvering to LVLH (yaw 0, pitch 0, roll 0) at 2:16pm in order for TsUP to perform additional data collection.  Normally, these tests do not call for maneuvering, but today’s testing was required to occur inthe same attitude the station will be in for the approach of the European ATV (automated transfer vehicle). As part of the test, MCC-H positioned the U.S. solar arrays to characterize multipath interference. Additional testing is taking place throughout the day without attitude maneuver or solar array position requirements.  Attitude control will return to U.S. CMG Momentum Management at 4:21pm.]

Update on ISS mass determination:   Using the data obtained from the reboost maneuver on 8/19, RSC-Energia is in the process of calculating the current mass of the ISS.  The effort, based on Newton’s law (force is equal to mass multiplied by acceleration), is estimated to take about two months, according to RSC-E.  The current mass estimate is 183 tons.  The calculation is expected to yield the exact value within 2% accuracy.

Major upcoming events:

  • ISS Reboost2 — 8/25, 8:32pm (8 min, delta-V = 2.2 m/s);
  • EVA-11 — 9/3;
  • Soyuz 9S launch — 10/9;
  • Soyuz 9S dock — 10/11;
  • Soyuz 8S undock/land — 10/19;
  • Soyuz 9S relocate — 11/18;
  • Progress 16P launch — 11/24.

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO photo targets, in the current LVLH attitude no longer limited by Flight Rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours) were Karachi, Pakistan (there was clear weather over the port city of Karachi.  This nadir pass provided an opportunity to capture the majority of the metropolitan area), Moroccan Dust Storm (Dynamic event.  A well-developed dust storm is active off of the northern coast of Morocco.  Looking to the left of track towards the southern coast of Spain), and Puerto Rico (the pass took ISS over the western edge of Puerto Rico.  Researchers have good reef data over this part of the island, but lack similar data for the eastern reefs.  Looking towards the eastern end of the island, as reefs here should be clearly visible.  Differences in magnitude of coral bleaching near urbanized areas are of particular interest).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:28pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 738; temperature (deg C) — 25.2; ppO2 (mmHg) — 161.4; ppCO2 (mmHg) — 3.4.
  • SM Transfer Compartment:  Pressure (mmHg) — 736; temperature (deg C) — 19.5.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 736.2; temperature (deg C) — 22.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 737.0; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 738.1; temperature (deg C) — 22.5; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg angle).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off line.  All other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System (PS):

  • Total propellant load available: 4432 kg (9771 lb) as of 8/19/04;  [SM(552) + FGB(3516) + Progress M(441)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7.2 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:39am EDT [= epoch]):

  • Mean altitude — 357.6 km
  • Apogee height — 363.0 km
  • Perigee height — 352.2 km
  • Period — 91.7 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0007978
  • Solar Beta Angle — 3.0 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32904

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.